OpenStudy (anonymous):

Ill fan and give medal! I need help making up 8 questions for the following Ellipses, hyperbolas, circles, parabolas radicals & exponents logs Rational expressions

OpenStudy (mathmale):

Please share with us an example of what you mean by "making up questions." Have you any study materials? Textbook? Online explanations? If you were to do an Internet search for "ellipse," you'd almost surely find examples of equations ellipses in the search results.

OpenStudy (anonymous):

nope. I just need to come up with 8 challenging questions of the following; Ellipses, hyperbolas, circles, parabolas radicals & exponents logs Rational expressions . id prefer from radicals down. i don't really want to do do anything having to do with Hyperbolas, circles, parabolas and ellipses.

OpenStudy (mathmale):

Looks like you'll need to invest some time and effort into researching these topics, at lest far enough to find some examples. If you're on the Internet, I'm sure you've done Internet searches before. Why not now? An example of an equation of an ellipse is\[\frac{ (x-1)^2 }{ 5^2 }+\frac{ (y+2)^2 }{ 3^2 }=1\] Look up the other terms for examples of equations of other conic sections.

OpenStudy (anonymous):

no idea BUT you ca give me a medal ;)

OpenStudy (mathmale):

Looks like you'll need to invest some time and effort into researching these topics, at lest far enough to find some examples. If you're on the Internet, I'm sure you've done Internet searches before. Why not now? Medals are awarded for good work (especially when the student makes effort to follow instructions).